OpenStudy (anonymous):

rationalize denominator: 5/8+√x

jimthompson5910 (jim_thompson5910):

\[\large \frac{5}{8+\sqrt{x}}\] \[\large \frac{5}{8+\sqrt{x}}\times\frac{8-\sqrt{x}}{8-\sqrt{x}}\] \[\large \frac{5(8-\sqrt{x})}{(8+\sqrt{x})(8-\sqrt{x})}\] \[\large \frac{40-5\sqrt{x}}{64-x}\] So \[\large \frac{5}{8+\sqrt{x}}=\frac{40-5\sqrt{x}}{64-x}\]

OpenStudy (anonymous):

is this fully simplified?

jimthompson5910 (jim_thompson5910):

Yes it's fully simplified

OpenStudy (anonymous):

dang, it's not correct...

OpenStudy (anonymous):

oh crap... sorry, it is right :)

jimthompson5910 (jim_thompson5910):

Yeah the computer is real picky